A family consists of a father, mother and some children. The mean age of the family is 18 years. Without the 38 year old father the mean age of the family decreases to only 14. How many children are there in the family?
Answer by oscargut(2103)

Let take n=number of children
h= sum of the ages of mother and children
then
(38+h)/(n+2) = 18
h/n+1 = 14
then
h=14(n+1)
then
38+14(n+1)=18(n+2)
then
38+14+14n=18n+36
-4n=-16
then
n=4
R: there are 4 children in the family